# Energy Dynamics Powered by Traction and Stress Control Formation and Motion of +1/2 Topological Defects in Epithelial Cell Monolayers

### v1

@@ Methods
with $\bar{P}=P-P_{m}$ and $P_{m}$ being the average of $P$ over space, $\bm{r}$ and $\bm{r}^{\prime}$ indicating positions in space, and the scalar $r$ being the magnitude of vector $\bm{r}$. The sum $\sum$ was taken over all positions in the cell island. The correlation length was defined as the distance over which $C(r)$ decreased to 0.
For analyzing confocal images, the upper (apical) $z$-planes of actin images were used to identify the cell boundaries using Cellpose version 3.0.7 for segmentation.39 The cell boundaries were fed to OrientationJ using a window size of 64 pix (19.84 µm) and the “Cubic Spline” option for computing the gradient. The orientation fields were overlaid with the cell boundaries to confirm accuracy of the segmentation (Supplemental Fig. S10d). Then the lower (basal) $z$-planes of the confocal actin image stacks were overlaid with the cell boundaries and the +1/2 defect to quantify manually the average orientation of stress fibers for each fixed cell on the defect axis (Supplemental Fig. S10e).
Statistical
@@ Acknowledgments
We thank Christian Franck for use of the confocal microscope. We thank Fridtjof Brauns for insightful discussions held at the Kavli Institute for Theoretical Physics (KITP), which is supported by NSF grant number PHY-2309135 and Gordon and Betty Moore Foundation Grant number 2919.02. This work was supported by the University of Wisconsin-Madison Office of the Vice Chancellor for Research and Graduate Education with funding from the Wisconsin Alumni Research Foundation, National Science Foundation grant number CMMI-2205141, and National Institutes of Health grant number 1R35GM151171.
@@
Energy Dynamics Powered by Traction and Stress Control Formation and Motion of +1/2 Topological Defects in Epithelial Cell Monolayers
Pradip K. Bera1∗, Molly McCord1,2∗, Jun Zhang1,3, Jacob Notbohm1,2,3
1 Department of Mechanical Engineering
2 Biophysics Program
3 Department of Nuclear Engineering and Engineering Physics
University of Wisconsin–Madison, Madison, Wisconsin 53706, USA
These authors contributed equally.
@@ 3 Department of Nuclear Engineering and Engineering Physics
Supplemental Figure S2. Additional examples of velocity field near +1/2 defects moving in the tail-to-head and head-to-tail directions for cells treated with CN02 (a) and CN03 (b). Yellow lines show cell orientations, white lines show velocity, and colors show vorticity.
Supplemental Figure S3. Plots of strain rate $\dot{\varepsilon}_{xx}$, stress $\sigma_{xx}$, and traction for a single head-to-tail defect and a single tail-to-head defect. Time points shown are 1 hr before defect formation ($t=-1$ hr) and immediately after defect formation ($t=0$).
@@
Supplemental Figure S5. Gradient of stress along the tail of +1/2 defects. (a,b) For head-to-tail moving defects, the average of $\sigma_{xx}$ is plotted against $x$ position along the defect tail (i.e., for $y=0$) at $t=-1$ hr (a) and at $t=0$ (b). (c,d) Similar plots for tail-to-head moving defects. Near the defect head (e.g., $0\leq x\leq 50$ µm), the gradient of $\sigma_{xx}$ is positive for both cases.
@@
Supplemental Figure S6. The gradient of stress and the average principal stress near +1/2 defects. (a,b) Considering head-to-tail defects, heat maps of number distribution in the [$<\!v_{x}\!>$, $<\!\partial\sigma_{xx}/\partial x\!>$] plane at $t=-1$ hr and at $t=0$ hr, respectively. Data were averaged along each defect tail ($0\leq x\leq 50$ µm, $y=0$). The plots indicate notable heterogeneity with the magnitude of $<\!v_{x}\!>$ varying from 0 to 20 µm/min and the magnitude of $<\!\partial\sigma_{xx}/\partial x\!>$ ranging from 0 to 2 Pa/µm. (c,d) For head-to-tail defects, map of average principal stress $(\sigma_{1}+\sigma_{2})/2$, and its variation along the tail (i.e., along $y=0$) at $t=0$ hr. (e-h) Similar plots for tail-to-head moving defects.
@@
Supplemental Figure S7. The gradient of stress via the divergence near +1/2 defects. (a,b) Considering head-to-tail defects, heat maps of number distribution in the [$<\!v_{x}\!>$, $<\!\partial\sigma_{xx}/\partial x+\partial\sigma_{xy}/\partial y\!>$] plane at $t=-1$ hr and at $t=0$ hr respectively. Data were averaged along each defect tail ($0\leq x\leq 50$ µm, $y=0$). Here, $\partial\sigma_{xx}/\partial x+\partial\sigma_{xy}/\partial y$ is the $x$ component of the stress equilibrium expression $\nabla\cdot\bm{\sigma}$, with $\bm{\sigma}$ being the stress tensor. (c) For head-to-tail defects, the number distribution heat map in the $<\!v_{x}\!>$ and $<\!\partial[(\sigma_{1}+\sigma_{2})/2]/\partial x\!>$ plane is shown at $t=0$ hr. (d-f) Similar plots for tail-to-head moving defects.
@@
Supplemental Figure S8. Angle $\theta$ between traction and velocity near +1/2 defects, and near defect-free counter rotating vortices. (a) The schematic illustrations show possible angles ($\theta$) between traction (applied by the substrate onto the cells) and cell velocity. If $\theta>90^{\circ}$ (red shading), traction acts as a friction; if $\theta<90^{\circ}$ (blue shading), traction is propulsive. (b) Histograms of $\theta$ at $t=0$ hr, considering all cells in 27 islands (gray) and cells located on the defect tail in the range $0\leq x\leq 100$ µm for head-to-tail (red) and tail-to-head (blue) defects, respectively. (c) Angle $\theta$ was also measured in 25 defect-free regions (length 100 µm, width 20 µm) having counter-rotating vortices. The distribution (orange) is nearly flat, similar to the histogram of $\theta$ for all cells (gray), which rules out the alternate explanation that counter rotating vortices themselves caused the tendency toward 0 and 180.
Supplemental Figure S9. Stress power density and traction power density near +1/2 defects at $t=-1$ hr and at $t=0$ hr for head-to-tail defects (a-d) and tail-to-head defects (e-h). The average power density associated with the stress and the traction are calculated according to $P_{S}=\bm{\sigma}:\dot{\bm{\varepsilon}}=\sigma_{xx}\dot{\varepsilon}_{xx}+2\sigma_{xy}\dot{\varepsilon}_{xy}+\sigma_{yy}\dot{\varepsilon}_{yy}$ and $P_{T}=-\bm{t}\cdot\bm{v}/h=-(t_{x}v_{x}+t_{y}v_{y})/h$, respectively, where $h$ is the cell height.
@@
Supplemental Figure S10. Figure showing how stress fiber orientations near +1/2 defects were identified. (a, b) Phase contrast image and image of fluorescent nuclei near a representative +1/2 defect before fixing cells. Yellow lines indicate cell orientation. The identified defect is indicated in red. (c) Confocal image of nuclei after fixing cells, which was used to identify the location of the defect of interest. (d) Confocal image of actin at the the apical (top) side of teh cell layer. White lines show segmented cell boundaries. Cell orientations were determined by applying OrientationJ to the the images of apical actin (shown in magenta) to verify the presence of a +1/2 defect. (e) Segmented cell boundaries from the apical images (shown in panel d) are overlaid on an image of the actin stress fibers on the basal (bottom) side of the cell monolayer. To quantify angle $\beta$, the orientations of stress fibers were determined manually for each cell along the defect tail. Scale bars represent 50 µm.
Video Captions
@@
Video 2. Zoomed version of the red region in Video 1. The red dot indicates the head of the head-to-tail moving +1/2 defect. The arrow on the defect tail indicates the movement of the defect.
Video 3. Zoomed version of the blue region in Video 1. The blue dot indicates the head of the tail-to-head moving +1/2 defect. The arrow on the defect tail indicates the movement of the defect.

### v2

@@ Methods
with $\bar{P}=P-P_{m}$ and $P_{m}$ being the average of $P$ over space, $\bm{r}$ and $\bm{r}^{\prime}$ indicating positions in space, and the scalar $r$ being the magnitude of vector $\bm{r}$. The sum $\sum$ was taken over all positions in the cell island. The correlation length was defined as the distance over which $C(r)$ decreased to 0.2.
Analysis of Confocal Images of Actin
For analyzing confocal images, the upper (apical) $z$-planes of actin images were used to identify the cell boundaries using Cellpose version 3.0.7 for segmentation.46 The cell boundaries were fed to OrientationJ using a window size of 64 pix (19.84 µm) and the “Cubic Spline” option for computing the gradient. The orientation fields were overlaid with the cell boundaries to confirm accuracy of the segmentation (Supplemental Fig. S10d). Then the lower (basal) $z$-planes of the confocal actin image stacks were overlaid with the cell boundaries and the +1/2 defect to quantify manually the average orientation of stress fibers for each fixed cell on the defect axis (Supplemental Fig. S10e).
Statistical Analysis
@@
Data and Code Availability
A partial data set of this study, including unprocessed images of cells and fluorescent particles, and processed data, is available at https://doi.org/10.6084/m9.figshare.28774310 (ref. 47). The full data set will be shared by the lead contact upon request. Code used to compute cell velocities, tractions, and stresses is available from https://github.com/jknotbohm/FIDIC and https://github.com/jknotbohm/Cell-Traction-Stress.
Author Contributions
P.K.B., M.M., J.Z., and J.N. designed the experiments. M.M. and J.Z. performed the experiments. P.K.B., J.Z., and J.N. analyzed data. P.K.B. and J.N. wrote the manuscript.
Declaration of Interests
The authors declare no competing interests.
@@ Acknowledgments
We thank Christian Franck for use of the confocal microscope. We thank Fridtjof Brauns for insightful discussions held at the Kavli Institute for Theoretical Physics (KITP), which is supported by NSF grant number PHY-2309135 and Gordon and Betty Moore Foundation Grant number 2919.02. This work was supported by the University of Wisconsin-Madison Office of the Vice Chancellor for Research and Graduate Education with funding from the Wisconsin Alumni Research Foundation, National Science Foundation grant number CMMI-2205141, and National Institutes of Health grant number R35GM151171.
@@
Supplemental Note 1: Derivation of Clapeyron’s Theorem
Here, we show the derivation for Clapeyron’s Theorem, which represents the balance between the three terms plotted in Fig. 4. We begin by integrating $\bm{\sigma}:\bm{\dot{\varepsilon}}$ over some region of interest $\Omega$, noting that the strain rate $\bm{\dot{\varepsilon}}$ is related to the gradient of velocity $\bm{v}$, enabling us to write
$\int_{\Omega}\bm{\sigma}:\bm{\dot{\varepsilon}}dA=\int_{\Omega}\bm{\sigma}:\frac{1}{2}\left[\nabla\bm{v}+(\nabla\bm{v})^{T}\right]dA=\int_{\Omega}\bm{\sigma}:(\nabla\bm{v})dA,$ (1)
where the last expression results from symmetry of the stress tensor. The right hand side of Eq. 1 can be rewritten to give
$\int_{\Omega}\bm{\sigma}:\bm{\dot{\varepsilon}}dA=\int_{\Omega}\nabla\cdot(\bm{\sigma}\bm{v})dA-\int_{\Omega}(\nabla\cdot\bm{\sigma})\cdot\bm{v}dA.$ (2)
The right hand side of Eq. 2 can be simplified by using the divergence theorem and noting that, by equilibrium, $\nabla\cdot\bm{\sigma}=-\bm{t}/h$ to give
$\int_{\Omega}\bm{\sigma}:\bm{\dot{\varepsilon}}dA=\int_{\Gamma}\bm{\sigma}\hat{\bm{n}}\cdot\bm{v}ds+\int_{\Omega}\frac{\bm{t}}{h}\cdot\bm{v}dA,$ (3)
where $\Gamma$ is the path enclosing region $\Omega$, and $\hat{\bm{n}}$ is the unit outward normal vector from $\Gamma$. Note that no constitutive relationship is assumed in deriving Eq. 3. This equation is Clapeyron’s Theorem, and it describes the flow of power into and out of the region of interest, with the first term on the right hand side representing power flowing through the boundaries and transmitted through stresses and the second term representing power generated or dissipated by tractions at the cell-substrate interface. Finally, as described in the main text, we define $P_{S}=\bm{\sigma}:\bm{\dot{\varepsilon}}$ and $P_{T}=-\bm{t}\cdot\bm{v}/h$, enabling us to rewrite Clapeyron’s Theorem as
$\int_{\Omega}P_{S}dA+\int_{\Omega}P_{T}dA=\int_{\Gamma}\bm{\sigma}\hat{\bm{n}}\cdot\bm{v}ds.$ (4)
The units of each integral above are the power per length, where the length is the height of the cell layer. Typically, the integrals above are each described as of a power; the reason for the discrepancy is that the integrals are written using a two-dimensional formulation, whereas our manuscript reports the stress $\bm{\sigma}$ as a three-dimensional quantity with units of force per length squared. To report values of power in this manuscript, we multiplied each integral by the height of the cell layer, $h$.
Supplemental Note 2: Estimation of Noise Floor in Power Density Measurements
Given that the powers $E_{S}$, $E_{T}$, and $E_{\Gamma}$ do not balance perfectly in Fig. 4j,k, here we analyze in detail the noise in the measurements of power. We consider the full cell island, for which $\bm{\sigma}\hat{\bm{n}}=0$, meaning $E_{\Gamma}=0$ and then we expect $E_{S}+E_{T}=0$. For this analysis, we choose to use units of power density (power/volume), as any error in the measurement can be directly compared to the values of power density shown in Fig. 4a,b. The quantities $E_{S}$ and $E_{T}$, calculated for the full cell island and normalized by the island volume, are shown in Fig. S11. As can be seen, their sum is in the range of 0.01–0.03 mW/m3, which is nearly a factor of 100 smaller than the typical values of power density shown in Fig. 4a,b, which are on the order of 1 mW/m3. Hence, the powers balance, and the level of noise in the data is negligible.
@@
Supplemental Figure S2. Additional examples of velocity field near +1/2 defects. Defects moving in the tail-to-head and head-to-tail directions for cells treated with CN02 (a) and CN03 (b). Yellow lines show cell orientations, white lines show velocity, and colors show vorticity.
Supplemental Figure S3. Plots of strain rate $\dot{\varepsilon}_{xx}$, stress $\sigma_{xx}$, and traction for a single head-to-tail defect and a single tail-to-head defect. Time points shown are 1 hr before defect formation (t=1t=-1 hr) and immediately after defect formation ($t=0$).
@@
Supplemental Figure S5. Gradient of stress along the tail of +1/2 defects. (a,b) For head-to-tail moving defects, the average of $\sigma_{xx}$ is plotted against $x$ position along the defect tail (i.e., for $y=0$) at t=1t=-1 hr (a) and at $t=0$ (b). (c,d) Similar plots for tail-to-head moving defects. Near the defect head (e.g., $0\leq x\leq 50$ µm), the gradient of $\sigma_{xx}$ is positive for both cases.
@@
Supplemental Figure S6. The gradient of stress and the average principal stress near +1/2 defects. (a,b) Considering head-to-tail defects, heat maps of number distribution in the $v_{x}\!>$, $\partial\sigma_{xx}/\partial x\!>$] plane at t=1t=-1 hr and at $t=0$ hr, respectively. Data were averaged along each defect tail ($0\leq x\leq 50$ µm, $y=0$). The plots indicate notable heterogeneity with the magnitude of $v_{x}\!>$ varying from 0 to 20 µm/min and the magnitude of $\partial\sigma_{xx}/\partial x\!>$ ranging from 0 to 2 Pa/µm. (c,d) For head-to-tail defects, map of average principal stress $(\sigma_{1}+\sigma_{2})/2$, and its variation along the tail (i.e., along $y=0$) at $t=0$ hr. (e-h) Similar plots for tail-to-head moving defects.
@@
Supplemental Figure S7. The gradient of stress via the divergence near +1/2 defects. (a,b) Considering head-to-tail defects, heat maps of number distribution in the $v_{x}\!>$, $\partial\sigma_{xx}/\partial x+\partial\sigma_{xy}/\partial y\!>$] plane at t=1t=-1 hr and at $t=0$ hr respectively. Data were averaged along each defect tail ($0\leq x\leq 50$ µm, $y=0$). Here, $\partial\sigma_{xx}/\partial x+\partial\sigma_{xy}/\partial y$ is the $x$ component of the stress equilibrium expression $\nabla\cdot\bm{\sigma}$, with $\bm{\sigma}$ being the stress tensor. (c) For head-to-tail defects, the number distribution heat map in the $v_{x}\!>$ and $\partial[(\sigma_{1}+\sigma_{2})/2]/\partial x\!>$ plane is shown at $t=0$ hr. (d-f) Similar plots for tail-to-head moving defects.
@@
Supplemental Figure S8. Angle $\theta$ between traction and velocity near +1/2 defects, and near defect-free counter-rotating vortices. (a) The schematic illustrations show possible angles ($\theta$) between traction (applied by the substrate onto the cells) and cell velocity. If $\theta>90^{\circ}$ (red shading), traction acts as a friction; if $\theta<90^{\circ}$ (blue shading), traction is propulsive. (b) Histograms of $\theta$ at $t=0$ hr, considering all cells in 27 islands (gray) and cells located on the defect tail in the range $0\leq x\leq 100$ µm for head-to-tail (red) and tail-to-head (blue) defects, respectively. (c) Angle $\theta$ was also measured in 25 defect-free regions (length 100 µm, width 20 µm) having counter-rotating vortices. The distribution (orange) is nearly flat, similar to the histogram of $\theta$ for all cells (gray), which rules out the alternate explanation that counter-rotating vortices themselves caused the tendency toward 0 and 180.
Supplemental Figure S9. Power flow at t=1t=-1 hr near +1/2 defects. (a, b, c) At t=1t=-1 hr near head-to-tail defects, power densities $P_{S}$, $P_{T}$, and the flow of power from the side cells $E_{\gamma}$ are shown for the ROI. (d–f) Similar plots are shown for tail-to-head moving defects at t=1t=-1 hr.
@@
Supplemental Figure S10. Figure showing how stress fiber orientations near +1/2 defects were identified. (a, b) Phase contrast image and image of fluorescent nuclei near a representative +1/2 defect before fixing cells. Yellow lines indicate cell orientation. The identified defect is indicated in red. (c) Confocal image of nuclei after fixing cells, which was used to identify the location of the defect of interest. (d) Confocal image of actin at the apical (top) side of the cell layer. White lines show segmented cell boundaries. Cell orientations were determined by applying OrientationJ to the the images of apical actin (shown in magenta) to verify the presence of a +1/2 defect. (e) Segmented cell boundaries from the apical images (shown in panel d) are overlaid on an image of the actin stress fibers on the basal (bottom) side of the cell monolayer. To quantify angle $\beta$, the orientations of stress fibers were determined manually for each cell along the defect tail. Scale bars represent 50 µm.
Supplemental Figure S11. Verification of global energy balance for a representative island. For a full cell island, $\bm{\sigma}\hat{\bm{n}}=0$, meaning $E_{\Gamma}=0$. (a) Power densities $E_{S}/V$ and $E_{T}/V$ are plotted over time for a representative full island, where $V$ is the volume of the cell island. (b) The sum $(E_{S}+E_{T})/V$ is shown over time. Differences from zero are due to experimental noise, which is small compared to the values power density shown in Fig. 4a,b.
Supplemental Video Captions
@@
Video 2. Zoomed in video of the red region in Video 1. The red dot indicates the head of the head-to-tail moving +1/2 defect. The arrow on the defect tail indicates the movement of the defect.
Video 3. Zoomed in video of the blue region in Video 1. The blue dot indicates the head of the tail-to-head moving +1/2 defect. The arrow on the defect tail indicates the movement of the defect.